# Accelerated Parallel Tempering via Neural Transports

### v2

@@ Abstract
Markov Chain Monte Carlo (MCMC) algorithms are essential tools in computational statistics for sampling from unnormalised probability distributions, but can be fragile when targeting high-dimensional, multimodal, or complex target distributions. Parallel Tempering (PT) enhances MCMC’s sample efficiency through annealing and parallel computation, propagating samples from tractable reference distributions to intractable targets via state swapping across interpolating distributions. The effectiveness of PT is limited by the often minimal overlap between adjacent distributions in challenging problems, which requires increasing the computational resources to compensate. We introduce a framework that accelerates PT by leveraging neural samplers—including normalising flows, diffusion models, and controlled diffusions—to reduce the required overlap. Our approach utilises neural samplers in parallel, circumventing the computational burden of neural samplers while preserving the asymptotic consistency of classical PT. We demonstrate theoretically and empirically on a variety of multimodal sampling problems that our method improves sample quality, reduces the computational cost compared to classical PT, and enables efficient free energies/normalising constants estimation.
@@ 1 Introduction
Sampling from a probability distribution $\pi(x)=\exp(-U(x))/Z$ defined over a state-space $\mathcal{X}$ with tractable potential $U:\mathcal{X}\to\mathbb{R}$ and intractable normalising constant $Z=\int_{\mathcal{X}}\exp(-U(x))\mathrm{d}x$ is a fundamental task in machine learning and natural sciences. Markov Chain Monte Carlo (MCMC) methods are usually employed for such purposes, constructing an ergodic Markov chain $(X_{t})_{t\in\mathbb{N}}$ using local moves leaving $\pi$ invariant. While MCMC algorithms are guaranteed to converge asymptotically, in practice, they struggle when the target is complex with multiple well-separated modes (Papamarkou et al., 2022; Hénin et al., 2022). To handle such cases, Parallel Tempering (PT) (Swendsen and Wang, 1986; Geyer, 1991; Hukushima and Nemoto, 1996) is a popular class of MCMC methods designed to improve the global mixing of locally efficient MCMC algorithms.
@@
PT works by considering an annealing path $\pi^{0},\pi^{1},\dots,\pi^{N}$ of distributions over $\mathcal{X}$ interpolating between a simple reference distribution $\pi^{0}=\eta$ (e.g. a Gaussian) and the target $\pi^{N}=\pi$. PT algorithms construct a Markov chain $\textbf{X}_{t}=(X_{t}^{0},\ldots,X_{t}^{N})$ on the extended state-space $\mathcal{X}^{N+1}$, targeting the joint distribution $\bm{\pi}=\pi^{0}\otimes\cdots\otimes\pi^{N}$. The PT chain $\mathbf{X}_{t}$ is constructed by alternating between (1) a local exploration phase where the $n$-th chain  [^1] Following the PT literature, we also refer to components of $\mathbf{X}_{t}$ as chains. of $\mathbf{X}_{t}$ is updated according to a $\pi^{n}$-invariant MCMC algorithm; and (2) a communication phase which proposes a sequence of swaps between neighbouring states accepted according to a Metropolis–Hastings correction ensuring invariance (see Figure 1 (a)). Crucially, PT offsets the additional computation burden of simulating the extended $N$ chains through parallel computation, allowing for a similar effective computational cost as a single chain when implemented in a maximally parallelised manner.
Figure 1: (Left) An illustration of the local exploration and communication step for PT vs APT. (Middle) 1,000 samples of a Gaussian mixture model target obtained using PT vs APT with a standard Gaussian reference. See Section 6.1 for more details. (Right) Round trips for PT and APT with $N=6$ chains over $T=100,000$ iterations of Algorithm 1.
@@
Typically, the chains $X_{t}^{n}$ mix faster when closer to the reference and struggle closer towards the target. Therefore, communication between the reference and target, facilitated through swaps, can induce rapid mixing between modes of the target component (Woodard et al., 2009; Surjanovic et al., 2024). While the importance of the swapping mechanism for PT has led to a literature dedicated to optimising communication between the reference and target (Syed et al., 2021, 2022; Surjanovic et al., 2022) such works still rely on the original swapping mechanism (Geyer, 1991).
@@
As an alternative to PT, recent work has explored both continuous (Zhang and Chen, 2022; Vargas et al., 2023; Berner et al., 2024; Akhound-Sadegh et al., 2024; Vargas et al., 2024; Máté and Fleuret, 2023; Albergo and Vanden-Eijnden, 2024; Erives et al., 2025) and discrete (Noé et al., 2019; Papamakarios et al., 2021; Midgley et al., 2023; Gabrié et al., 2022) flows for sampling, under the umbrella of neural samplers. However, these methods usually incur a bias, foregoing theoretical guarantees of MCMC, and can be expensive to implement and train. Due to these shortcomings, standard PT provides a strong baseline that neural samplers struggle to match (He et al., 2025b).
@@
Recent work (Arbel et al., 2021; Albergo and Vanden-Eijnden, 2024; Phillips et al., 2024; Chen et al., 2025) has explored approaches to debias neural samplers using Sequential Monte Carlo (SMC)-based ideas (Del Moral et al., 2006). Despite their consistency guarantees, these approaches do not address the mode-collapsing nature of many modern neural samplers (He et al., 2025b). PT offers a comparable but computationally dual framework to SMC, where the roles of parallelism and time are reversed (Syed et al., 2024). In SMC, particles are generated in parallel, and approximate annealing distributions are constructed sequentially. In contrast, in PT, particles are generated sequentially and annealing distributions are built in parallel. This raises the question: just as neural samplers have been integrated with SMC, can we integrate neural samplers with PT, combining the consistency of PT with the flexibility of neural samplers?
@@
We answer positively this question by formalising and exploiting the framework introduced by Ballard and Jarzynski (2009, 2012) in physics for designing more flexible swap mechanisms, which we call accelerated PT (APT). APT preserves PT’s asymptotic consistency and allows us to integrate easily normalising flows, stochastic control, and diffusions into existing PT implementations. ATP uses neural sampling in a parallelised manner, mitigating its high computational burden. Empirically, it outperforms PT by accelerating the communication between the reference and target states, even when controlling for the additional computation incurred by neural samplers.
@@ 2 Parallel Tempering
There is considerable flexibility in choosing annealing distributions provided $U^{0}(x)=\log\eta(x)$ with $Z_{0}=1$ and $U^{N}(x)=U(x)$ with $Z_{N}=Z$. Without loss of generality, we can assume $U^{n}=U^{\beta_{n}}$ where for $\beta\in[0,1]$, $U^{\beta}:\mathcal{X}\to\mathbb{R}$ continuously interpolates between reference and target potentials as $\beta$ increases from 00 to $1$, according to the annealing schedule $10=\beta_{0}<\cdots<\beta_{N}=$ 1. A common choice is the linear path, $U^{\beta}(x)=(1-\beta)\log\eta(x)+\beta U(x)$, which linearly interpolates between reference and target in log-space. See Masrani et al. (2021); Syed et al. (2021); Máté and Fleuret (2023); York (2023) for alternative non-linear annealing paths.
@@ 2.1 Non-Reversible Parallel Tempering
The PT algorithm constructs a Markov chain $\mathbf{X}_{t}=(X_{t}^{0},\ldots,X_{t}^{N})$ on the extended state-space $\mathcal{X}^{N+1}$ invariant to the joint distribution $\bm{\pi}:=\pi^{0}\otimes\cdots\otimes\pi^{N}$. We construct $\mathbf{X}_{t}$ from $\mathbf{X}_{t-1}$ by doing (1) a local exploration step followed by (2) a communication step seen in the top of Figure 1 (a). For $n=0,\dots,N$, the $n$-th local exploration move (1) updates the $n$-th component of $\mathbf{X}_{t-1}$ using a $\pi^{n}$-invariant Markov kernel $K^{n}(x,\mathrm{d}x^{\prime})$ on $\mathcal{X}$ corresponding to an MCMC move targetting $\pi^{n}$,
@@
We additionally assume that $K^{0}(x,\mathrm{d}x^{\prime})=\eta(\mathrm{d}x^{\prime})$ corresponds to an independent sample from the reference. The communication step (2) applies a sequence of swap moves between adjacent components of $\mathbf{X}_{t}$, where the $n$-th swap move illustrated in Figure 1 exchanges components $X^{n-1}_{t}$ and $X^{n}_{t}$ in $\textbf{X}_{t}=(X^{0}_{t},\dots,X^{N}_{t})$ with probability $\alpha^{n}(X^{n-1}_{t},X^{n}_{t})$, where for $x,x^{\prime}\in\mathcal{X}$,
@@
While the effective sample size (ESS) of samples generated by a Markov chain is the gold standard for evaluating the performance of MCMC algorithms, in our setting, we are mainly interested in improving the swap kernel within PT. As ESS measures the intertwined performance of the local exploration and swap kernels, we are instead interested in maximising communication between reference and target, which can be empirically measured by counting the total number of round trips, $R_{T}$ tracking the number of independent reference samples transported to the target after $T$ iterations of PT (Katzgraber et al., 2006; Lingenheil et al., 2009). See Section A.1 for a formal definition. The mixing time of the PT chain is related to the time it takes for a round trip to occur and the total number of round trips is a measure of the particle diversity generated by PT and strongly correlates with the ESS (Surjanovic et al., 2024).
3 Accelerated Parallel Tempering
A limitation of PT is its inflexible swap move, which only proposes directly exchangeable samples between distributions, considering just the relative change in potentials $\pi^{n-1}$ and $\pi^{n}$. Low acceptance probability occurs when these distributions have minimal overlap. Addressing this requires increasing the number of parallel chains $N$, which may not always be possible. We propose Accelerated PT (APT), expanding the framework developed in Ballard and Jarzynski (2009, 2012) to improve distributional overlap and accelerate communication for a fixed number of chains.
3.1 Forward and Backward Accelerators
We assume ${\mathbb{P}}^{n-1}_{K}$ and ${\mathbb{Q}}^{n}_{K}$ are mutually absolutely continuous and we can evaluate $W^{n}_{K}:\mathcal{X}^{K+1}\to\mathbb{R}$, the work between the forward and the backward accelerated paths, extending $W^{n}$ in Equation 1,
$W^{n}_{K}(x_{0:K}):=\Delta F^{n}-\log\frac{\mathrm{d}{{\mathbb{Q}}}^{n}_{K}}{\mathrm{d}{{\mathbb{P}}}^{n-1}_{K}}(x_{0:K}).$ (3)
3.2 Non-Reversible Accelerated Parallel Tempering
We construct the Markov chain $\textbf{X}_{t}=(X^{0}_{t},\dots,X^{N}_{t})$ for $t=1,\dots,T$ using the same local exploration and non-reversible communication as classical PT, but we use using the accelerated PT swap move as described below and summarized in Algorithm 1. Given PT state $\textbf{X}_{t}=(X^{0}_{t},\dots,X^{N}_{t})\in\mathcal{X}^{N+1}$ after the local exploration move, we define the $n$-th accelerated swap move as follows: generate paths $\accentset{\rightharpoonup}{X}^{n-1}_{t,0:K}$, and $\accentset{\leftharpoonup}{X}^{n}_{t,0:K}$ obtained by propagating $X^{n-1}_{t}$ and $X^{n}_{t}$ forward and backward in time using the forward and backward transitions respectively,
The $n$-th accelerated swap move illustrated in Figure 1 replaces components $X^{n-1}_{t}$ and $X^{n}_{t}$ in $\textbf{X}_{t}$ with $\accentset{\leftharpoonup}{X}^{n}_{t,0}$ and $\accentset{\rightharpoonup}{X}^{n-1}_{t,K}$ respectively with probability $\alpha^{n}_{K}(\accentset{\rightharpoonup}{X}^{n-1}_{t,0:K},\accentset{\leftharpoonup}{X}^{n}_{t,0:K})$, where for $x_{0:K},x_{0:K}^{\prime}\in\mathcal{X}^{K+1}$,
When $K=0$, the accelerated swap coincides with the traditional PT swap in Equation 2. However, an accelerated swap can obtain an acceptance of $1$ even if $\pi^{n-1}\neq\pi^{n}$ provided ${\mathbb{P}}^{n-1}_{K}={\mathbb{Q}}^{n}_{K}$. Therefore, we aim to choose the forward and backward accelerators to make the laws of simulated forward and backward paths as close to each other as possible. Theorem 1 shows we can quantify this discrepancy in through the rejection rate, $r({\mathbb{P}}^{n-1}_{K},{\mathbb{Q}}^{n}_{K}):=\|{\mathbb{P}}^{n-1}_{K}\otimes{\mathbb{Q}}^{n}_{K}-{\mathbb{Q}}^{n}_{K}\otimes{\mathbb{P}}^{n-1}_{K}\|_{\mathrm{TV}}$ which by Pinsker inequality is controlled by the symmetric KL divergence between ${\mathbb{P}}^{n-1}_{K}$ and ${\mathbb{Q}}^{n}_{K}$,
@@ 3.2 Non-Reversible Accelerated Parallel Tempering
7:              $\accentset{\rightharpoonup}{X}^{n-1}_{t,k}\sim P_{k}^{n-1}(\accentset{\rightharpoonup}{X}^{n-1}_{t,{k-1}},\mathrm{d}x)$triangleright Accelerate forward
8:              $\accentset{\leftharpoonup}{X}^{n}_{t,K-k}\sim Q_{K-k}^{n}(\accentset{\leftharpoonup}{X}^{n}_{t,K-k+1},\mathrm{d}x)$triangleright Accelerate backward

### v3

@@ Abstract
Markov Chain Monte Carlo (MCMC) algorithms are essential tools in computational statistics for sampling from unnormalised probability distributions, but can be fragile when targeting high-dimensional, multimodal, or complex target distributions. Parallel Tempering (PT) enhances MCMC’s sample efficiency through annealing and parallel computation, propagating samples from tractable reference distributions to intractable targets via state swapping across interpolating distributions. The effectiveness of PT is limited by the often minimal overlap between adjacent distributions in challenging problems, which requires increasing the computational resources to compensate. We introduce a framework that accelerates PT by leveraging neural samplers—including normalising flows, diffusion models, and controlled diffusions—to reduce the required overlap. Our approach utilises neural samplers in parallel, circumventing the computational burden of neural samplers while preserving the asymptotic consistency of classical PT. We demonstrate theoretically and empirically on a variety of multimodal sampling problems that our method improves sample quality, reduces the computational cost compared to classical PT, and enables efficient free energy/normalising constant estimation.
@@ 1 Introduction
Sampling from a probability distribution $\pi(x)=\exp(-U(x))/Z$ defined over a state-space $\mathcal{X}$ with tractable potential $U:\mathcal{X}\to\mathbb{R}$ and intractable normalising constant $Z=\int_{\mathcal{X}}\exp(-U(x))\mathrm{d}x$ is a fundamental task in machine learning and natural sciences. Markov Chain Monte Carlo (MCMC) methods are usually employed for such purposes, constructing an ergodic Markov chain $(X_{t})_{t\in\mathbb{N}}$ using local moves leaving $\pi$ invariant. While MCMC algorithms are guaranteed to converge asymptotically, in practice, they struggle when the target is complex with multiple well-separated modes (Papamarkou et al., 2022; Hénin et al., 2022). To handle such cases, Parallel Tempering (PT) (Swendsen & Wang, 1986; Geyer, 1991; Hukushima & Nemoto, 1996) is a popular class of MCMC methods designed to improve the global mixing of locally efficient MCMC algorithms.
@@
PT works by considering an annealing path $\pi^{0},\pi^{1},\dots,\pi^{N}$ of distributions over $\mathcal{X}$ interpolating between a simple reference distribution $\pi^{0}=\eta$ (e.g. a Gaussian) and the target $\pi^{N}=\pi$. PT algorithms construct a Markov chain $\textbf{X}_{t}=(X_{t}^{0},\ldots,X_{t}^{N})$ on the extended state-space $\mathcal{X}^{N+1}$, targeting the joint distribution $\bm{\pi}=\pi^{0}\otimes\cdots\otimes\pi^{N}$. The PT chain $\mathbf{X}_{t}$ is constructed by alternating between (1) a local exploration phase where the $n$-th chain  [^1] Following the PT literature, we also refer to components of $\mathbf{X}_{t}$ as chains. of $\mathbf{X}_{t}$ is updated according to a $\pi^{n}$-invariant MCMC algorithm; and (2) a communication phase which proposes a sequence of swaps between neighbouring states accepted according to a Metropolis–Hastings correction ensuring invariance (see Figure˜1 (a)). Crucially, PT offsets the additional computation burden of simulating the extended $N$ chains through parallel computation, allowing for a similar effective computational cost as a single chain when implemented in a maximally parallelised manner.
@@
Typically, the chains $X_{t}^{n}$ mix faster when closer to the reference and struggle closer towards the target. Therefore, communication between the reference and target, facilitated through swaps, can induce rapid mixing between modes of the target component (Woodard et al., 2009; Surjanovic et al., 2024). While the importance of the swapping mechanism for PT has led to a literature dedicated to optimising communication between the reference and target (Syed et al., 2021; 2022; Surjanovic et al., 2022) such works still rely on the original swapping mechanism (Geyer, 1991).
Figure 1: (Left) An illustration of the local exploration and communication step for PT vs APT. (Middle) 1,000 samples of a Gaussian mixture model target obtained using PT vs APT with a standard Gaussian reference. See Section˜6.1 for more details. (Right) Round trips for PT and APT with $N=6$ chains over $T=100,000$ iterations of Algorithm˜1.
@@
As an alternative to PT, recent work has explored both continuous (Zhang & Chen, 2022; Vargas et al., 2023; Berner et al., 2024; Akhound-Sadegh et al., 2024; Vargas et al., 2024; Máté & Fleuret, 2023; Albergo & Vanden-Eijnden, 2024; Erives et al., 2025) and discrete (Noé et al., 2019; Papamakarios et al., 2021; Midgley et al., 2023; Gabrié et al., 2022) flows for sampling, under the umbrella of neural samplers. However, these methods usually incur a bias, foregoing theoretical guarantees of MCMC, and can be expensive to implement and train. Due to these shortcomings, standard PT provides a strong baseline that neural samplers struggle to match (He et al., 2025b).
@@
Recent work (Arbel et al., 2021; Albergo & Vanden-Eijnden, 2024; Phillips et al., 2024; Chen et al., 2025) has explored approaches to debias neural samplers using Sequential Monte Carlo (SMC)-based ideas (Del Moral et al., 2006). Despite their consistency guarantees, these approaches do not address the mode-collapsing nature of many modern neural samplers (He et al., 2025b). PT offers a comparable but computationally dual framework to SMC, where the roles of parallelism and time are reversed (Syed et al., 2024). In SMC, particles are generated in parallel, and approximate annealing distributions are constructed sequentially. In contrast, in PT, particles are generated sequentially and annealing distributions are built in parallel. This raises the question: just as neural samplers have been integrated with SMC, can we integrate neural samplers with PT, combining the consistency of PT with the flexibility of neural samplers?
@@
We answer positively this question by formalising and exploiting the framework introduced by Ballard & Jarzynski (2009; 2012) in physics for designing more flexible swap mechanisms, which we call accelerated PT (APT). APT preserves PT’s asymptotic consistency and allows us to integrate easily normalising flows, stochastic control, and diffusions into existing PT implementations. ATP uses neural sampling in a parallelised manner, mitigating its high computational burden. Empirically, it outperforms PT by accelerating the communication between the reference and target states, even when controlling for the additional computation incurred by neural samplers.
@@ 2 Parallel Tempering
There is considerable flexibility in choosing annealing distributions provided $U^{0}(x)=\log\eta(x)$ with $Z_{0}=1$ and $U^{N}(x)=U(x)$ with $Z_{N}=Z$. Without loss of generality, we can assume $U^{n}=U^{\beta_{n}}$ where for $\beta\in[0,1]$, $U^{\beta}:\mathcal{X}\to\mathbb{R}$ continuously interpolates between reference and target potentials as $\beta$ increases from 0 to $1$, according to the annealing schedule $0=\beta_{0}<\cdots<\beta_{N}=1$. A common choice is the linear path, $U^{\beta}(x)=(1-\beta)\log\eta(x)+\beta U(x)$, which linearly interpolates between reference and target in log-space. See Masrani et al. (2021); Syed et al. (2021); Máté & Fleuret (2023); York (2023) for alternative non-linear annealing paths.
@@ 2.1 Non-Reversible Parallel Tempering
The PT algorithm constructs a Markov chain $\mathbf{X}_{t}=(X_{t}^{0},\ldots,X_{t}^{N})$ on the extended state-space $\mathcal{X}^{N+1}$ invariant to the joint distribution $\bm{\pi}:=\pi^{0}\otimes\cdots\otimes\pi^{N}$. We construct $\mathbf{X}_{t}$ from $\mathbf{X}_{t-1}$ by doing (1) a local exploration step followed by (2) a communication step seen in the top of Figure˜1 (a). For $n=0,\dots,N$, the $n$-th local exploration move (1) updates the $n$-th component of $\mathbf{X}_{t-1}$ using a $\pi^{n}$-invariant Markov kernel $K^{n}(x,\mathrm{d}x^{\prime})$ on $\mathcal{X}$ corresponding to an MCMC move targetting $\pi^{n}$,
@@
We additionally assume that $K^{0}(x,\mathrm{d}x^{\prime})=\eta(\mathrm{d}x^{\prime})$ corresponds to an independent sample from the reference. The communication step (2) applies a sequence of swap moves between adjacent components of $\mathbf{X}_{t}$, where the $n$-th swap move illustrated in Figure˜1 exchanges components $X^{n-1}_{t}$ and $X^{n}_{t}$ in $\textbf{X}_{t}=(X^{0}_{t},\dots,X^{N}_{t})$ with probability $\alpha^{n}(X^{n-1}_{t},X^{n}_{t})$, where for $x,x^{\prime}\in\mathcal{X}$,
@@
While the effective sample size (ESS) of samples generated by a Markov chain is the gold standard for evaluating the performance of MCMC algorithms, in our setting, we are mainly interested in improving the swap kernel within PT. As ESS measures the intertwined performance of the local exploration and swap kernels, we are instead interested in maximising communication between reference and target, which can be empirically measured by counting the total number of round trips, $R_{T}$ tracking the number of independent reference samples transported to the target after $T$ iterations of PT (Katzgraber et al., 2006; Lingenheil et al., 2009). See Section˜A.1 for a formal definition. The mixing time of the PT chain is related to the time it takes for a round trip to occur and the total number of round trips is a measure of the particle diversity generated by PT and strongly correlates with the ESS (Surjanovic et al., 2024).
@@
7:     $\accentset{\rightharpoonup}{X}^{n-1}_{t,k}\sim P_{k}^{n-1}(\accentset{\rightharpoonup}{X}^{n-1}_{t,{k-1}},\mathrm{d}x)$ $\triangleright$ Accelerate forward
8:     $\accentset{\leftharpoonup}{X}^{n}_{t,K-k}\sim Q_{K-k}^{n}(\accentset{\leftharpoonup}{X}^{n}_{t,K-k+1},\mathrm{d}x)$ $\triangleright$ Accelerate backward
@@
3 Accelerated Parallel Tempering
A limitation of PT is its inflexible swap move, which only proposes directly exchangeable samples between distributions, considering just the relative change in potentials $\pi^{n-1}$ and $\pi^{n}$. Low acceptance probability occurs when these distributions have minimal overlap. Addressing this requires increasing the number of parallel chains $N$, which may not always be possible. We propose Accelerated PT (APT), expanding the framework developed in Ballard & Jarzynski (2009; 2012) to improve distributional overlap and accelerate communication for a fixed number of chains.
3.1 Forward and Backward Accelerators
We assume ${\mathbb{P}}^{n-1}_{K}$ and ${\mathbb{Q}}^{n}_{K}$ are mutually absolutely continuous and we can evaluate $W^{n}_{K}:\mathcal{X}^{K+1}\to\mathbb{R}$, the work between the forward and the backward accelerated paths, extending $W^{n}$ in Equation˜1,
$W^{n}_{K}(x_{0:K}):=\Delta F_{n}-\log\frac{\mathrm{d}{{\mathbb{Q}}}^{n}_{K}}{\mathrm{d}{{\mathbb{P}}}^{n-1}_{K}}(x_{0:K}).$ (3)
3.2 Non-Reversible Accelerated Parallel Tempering
We construct the Markov chain $\textbf{X}_{t}=(X^{0}_{t},\dots,X^{N}_{t})$ for $t=1,\dots,T$ using the same local exploration and non-reversible communication as classical PT, but we use using the accelerated PT swap move as described below and summarized in Algorithm˜1. Given PT state $\textbf{X}_{t}=(X^{0}_{t},\dots,X^{N}_{t})\in\mathcal{X}^{N+1}$ after the local exploration move, we define the $n$-th accelerated swap move as follows: generate paths $\accentset{\rightharpoonup}{X}^{n-1}_{t,0:K}$, and $\accentset{\leftharpoonup}{X}^{n}_{t,0:K}$ obtained by propagating $X^{n-1}_{t}$ and $X^{n}_{t}$ forward and backward in time using the forward and backward transitions respectively,
The $n$-th accelerated swap move illustrated in Figure˜1 replaces components $X^{n-1}_{t}$ and $X^{n}_{t}$ in $\textbf{X}_{t}$ with $\accentset{\leftharpoonup}{X}^{n}_{t,0}$ and $\accentset{\rightharpoonup}{X}^{n-1}_{t,K}$ respectively with probability $\alpha^{n}_{K}(\accentset{\rightharpoonup}{X}^{n-1}_{t,0:K},\accentset{\leftharpoonup}{X}^{n}_{t,0:K})$, where for $x_{0:K},x_{0:K}^{\prime}\in\mathcal{X}^{K+1}$,
When $K=0$, the accelerated swap coincides with the traditional PT swap in Equation˜2. However, an accelerated swap can obtain an acceptance of $1$ even if $\pi^{n-1}\neq\pi^{n}$ provided ${\mathbb{P}}^{n-1}_{K}={\mathbb{Q}}^{n}_{K}$. Therefore, we aim to choose the forward and backward accelerators to make the laws of simulated forward and backward paths as close to each other as possible. Theorem˜1 shows we can quantify this discrepancy in through the rejection rate, $r({\mathbb{P}}^{n-1}_{K},{\mathbb{Q}}^{n}_{K}):=\|{\mathbb{P}}^{n-1}_{K}\otimes{\mathbb{Q}}^{n}_{K}-{\mathbb{Q}}^{n}_{K}\otimes{\mathbb{P}}^{n-1}_{K}\|_{\mathrm{TV}}$ which by Pinsker inequality is controlled by the symmetric KL divergence between ${\mathbb{P}}^{n-1}_{K}$ and ${\mathbb{Q}}^{n}_{K}$,